Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 114; σ = 17 P(x ≥ 90) =
Answer #1

Given: 114 and u= = 17 When,x= 90 z .X-11 σ = -1.41 Therefore p(x 290 )= p (22 -1.41 ) = P(Z < 1.41 )= 0.9207 By standard nor